### v2

@@ 4.3 Results
Figure 1 presents point-wise relative bias ($\times$ 100) and 95% confidence interval coverage for the homogeneous and heterogeneous QTE scenarios. Throughout, we see that the proposed estimators (III) and (IV) in Section 4.2 have relative bias and confidence interval coverage comparable to the estimator (I) based on the full data. Regarding the 95% confidence bands for uniform inference, coverage was 96.7% for estimation of the homogeneous QTE and 98.0% for estimation of the heterogeneous QTE.
Finally, Figure 2 illustrates how results might be presented in practice, that is for a single dataset. In particular, Figure 2 compares 95% pointwise confidence intervals, including both the asymptotic variance formula-based and bootstrap-based intervals, versus the confidence bands for a realization, both for homogeneous and heterogeneous QTE settings. The black dotted line shows the point QTE estimates, the blue and yellow dashed lines show the 95% pointwise confidence intervals, and the light grey areas show the 95% uniform confidence bands, which are wider by construction. In the left panel of Figure 2, under a homogeneous QTE setting, the QTE estimates demonstrate similarity across various quantile levels; whereas in the right panel, the QTE estimates show more beneficial treatment effects with larger $Y$ values.
@@ 5 Data Application
In this section, we illustrate proposed methods proposed with a comparison of the relative effects of two bariatric surgery procedures on long-term weight loss using EHR data from a broader investigation of long-term outcomes for patients who had undergone bariatric surgery in one of the three Kaiser Permanente health care sites: Washington, Northern California, and Southern California(Arterburn et al., 2021). At each study site, information on health system enrollment, diagnoses, procedures, medical records, laboratory values, and other data systems were routinely recorded for research.
In the present study, we aim to compare Roux-en-Y gastric bypass (RYGB) ($Z=1$) versus vertical sleeve gastrectomy (VSG) ($Z=0$) bariatric surgery procedures on percent BMI change at 3 years post-surgery ($Y$). From a larger sample of 30,991 patients for which follow-up outcomes were partially observed, we selected $13,514$ patients who underwent RYGB or VSG between January 2005 and September 2015, with complete weight data at baseline (closest measurement pre-surgery, up to 6 months) and at follow-up (closest measurement within $\pm$ 90 days). See Table S.3 in the Supplementary Materials for a summary of the baseline characteristics of individuals who underwent bariatric surgery procedures.
@@
Subsequently, to impose MNAR, we additionally included a dependence on the outcome $Y$ by augmenting the fitted values $\widetilde{p}$ of $\textrm{pr}(R=1\mid Z,X)$ via:
[…]
where ${Y^{*}}$ is the standardized BMI change at 3 years. Finally, we trimmed $p$ using values $0.05$ and $0.95$.
To generate MNAR outcomes on the sample of 13,514 patients, we sampled $R$ according to a Bernoulli with probability given by the trimmed $p$. The resulting marginal probability of missingness was 30%. We considered double sampling random subsamples of size 1,000 and 1,500 from all patients with initially missing outcomes. For each of the two resulting data sets, we computed the QTEs at a sequence of quantile levels (i.e., $0.1,0.3,0.5,0.7,0.9$), and compared the point estimates and 95% pointwise bootstrap confidence intervals using the proposed method in Section 3. A benchmark QTE estimate for each quantile level was estimated using the full data set (before artificial imposition of missing outcomes). For comparison, we also estimated QTEs using the complete-case data set ($R=1$). We computed 95% confidence bands using the data set with 1,500 double-sampled patients and compared to the pointwise bootstrap confidence intervals. Additionally, as the effects of the two bariatric surgeries may be different among patient with diabetes or dyslipidemia diagnosis, we repeated the above analyses stratifying by diabetes and dyslipidemia diagnosis categories.
The left column of Figure 3 displays the pointwise QTE estimates and 95% confidence intervals across the quantile levels, while the right column compares pointwise CIs to confidence bands constructed using the procedure in Section 3.6. Panels (A) to (C) of Figure 3 represent different subsets of the data: (A) are for the overall 13,514 patients, while row (B) and (C) are stratified by diabetes status and dyslipidemia diagnosis, respectively. The specific estimated QTE values and 95% pointwise CIs are listed in Table S.4 in the Supplementary Materials.
From the results we see that the complete case estimates are biased when compared to the benchmark estimates that use the full data, especially at low quantile levels. The proposed IPW-WTQE estimators appeared to be unbiased, with the setting with 1,500 subjects in the second-stage sample showed greater efficiency (i.e. with narrower 95% CIs) compared to the setting with the smaller subsample. Heterogeneous QTEs were observed across quantile levels as well as at different strata: when compared to VSG, RYGB appears to be associated with greater percent BMI loss at 3 years: (i) at low tails of the percent BMI change distribution, and (ii) among patients with diabetes or dyslipidemia diagnosis at most of the selected quantile levels.
@@ 6 Discussion
This paper proposes a novel approach for estimation and inference for causal weighted QTEs in settings where outcome data are potentially MNAR. The approach combines additional data collection on a sub-sample of study units with otherwise missing information, with a set of proposed identifying assumptions. When estimating contrasts that involve the mean of the counterfactual distribution such as the ATE, one may employ $g$-methods that first compute a conditional ATE, conditional on $X$, and then marginalize to recover an estimate of the population causal effect (Robins, 1986; Hernán and Robins, 2020). However, due to Jensen’s inequality, integrating a conditional QTE will not yield the population QTE (Sun et al., 2021). In this work, therefore, we developed a novel inverse-probability weighted estimator as well as a bootstrap-based approach for valid inference pointwise for a single quantile or uniformly for a range of quantiles. The procedure involves first estimating propensity scores and double-sampling probabilities, and then plugging these into an estimating function for the weighted QTE. In a simulation study, we consider settings where the QTEs are either homogeneous or heterogeneous across quantiles, verify theoretical estimation properties in a finite sample setting and demonstrate validity of the proposed bootstrap-based confidence intervals/bands.
@@
Key to the proposed methods is that no assumptions are required regarding the missingness mechanism in the original data (i.e., for $R$). This, arguably, is a critical benefit in that the mechanism(s) underlying $R$—especially in the increasingly common setting in which EHR data are being used—are typically outside of the control of the research team and potentially quite complex. As such, even thought experiments regarding assumptions required to make progress with the original data (e.g., to consider the potential validity of MAR) can be challenging and inconclusive. One way of framing the proposed methods, therefore, is that the need to make assumptions regarding $R$ (as would have to be the case if only the original data were available) are replaced with the need to make assumptions regarding the double-sampling indicator $S$; that is, Assumptions 1 and 2. In practice, depending on the context and the nature of the additional data to be collected, consideration of these assumptions may be fairly straightforward. For example, suppose the additional data collection involves manual chart review or natural language processing of a sub-sample chosen by the research team (Weiskopf et al., 2019). In this setting, since no involvement of the study subject is needed, Assumptions 1 and 2 may hold trivially by design. In other settings, however, in particular when participation of a study subject is needed (e.g., to respond to a survey or phone call), assessment of these assumptions may require more detailed consideration. However, even in these situations, researchers may pursue additional design options that may shed light. For example, researchers may employ intensive, repeated efforts to contact study subjects and collect the data and/or engage with study subjects who have complete information in the original data to learn about potential recall bias (Haneuse et al., 2016; Koffman et al., 2021). EHR studies often encompass vast datasets, yet the ability to reach out to patients typically does not grow with sample size. Given that the double-sampling framework can be strategically planned from the beginning, it presents an opportunity to allocate resources effectively within the budgetary confines. Our ongoing research aims to find an optimal sampling rule that minimizes the semiparametric efficiency bound subject to a budget constraint.
@@
7 Supplementary Materials
Supplementary material available online includes the proofs of the results in the main manuscript, the gradient bootstrap computational algorithm, and additional tables and figures.
8 Competing interests
No competing interest is declared.
9 Acknowledgements

### v3

@@ 4.3 Results
Figure 1 presents point-wise relative bias ($\times$ 100) and 95% confidence interval coverage for the homogeneous and heterogeneous QTE scenarios. Throughout, we see that the proposed estimators (III) and (IV) in Section 4.2 have relative bias and confidence interval coverage comparable to the estimator (I) based on the full data. Regarding the 95% confidence bands for uniform inference, coverage was 98.6% for estimation of the homogeneous QTE and 98.9% for estimation of the heterogeneous QTE. Note, this over coverage is consistent with prior reports regarding confidence band estimation in finite samples (Belloni et al., 2019; Song et al., 2012; Kocherginsky et al., 2005). Finally, the complete case estimator (II) and estimator (V) assuming MAR are both substantially biased.
Figure 2 illustrates results for a single dataset, comparing 95% pointwise confidence intervals (asymptotic and bootstrap-based) with confidence bands in homogeneous and heterogeneous QTE settings. The black dotted line shows the point QTE estimates, the blue and yellow dashed lines show the 95% pointwise CIs, and the light grey areas show the 95% uniform confidence bands, which are wider by construction. In the left panel, QTE estimates demonstrate similarity across quantiles; whereas in the right panel, QTE estimates show more beneficial treatment effects with larger $Y$ values. Notably, the heavy-tailed error term $\varepsilon$ widens pointwise CIs on the right due to the density function in the variance formula (Theorem 3.1). Thus, the disparity between confidence bands and intervals narrows in the upper right of Figure 2. Similar patterns show with missingness rates of 20%, 50%, and 70% and smaller samples (2,000 and 5,000) (Section S.4.3 of the Supplementary Materials).
@@ 5 Data Application
In this section, we illustrate the proposed methods with a comparison of the relative effects of two bariatric surgery procedures on long-term weight loss using EHR data from a broader investigation of long-term outcomes for patients who had undergone bariatric surgery in one of the three Kaiser Permanente health care sites: Washington, Northern California, and Southern California(Arterburn et al., 2021). In the present study, we aim to compare Roux-en-Y gastric bypass (RYGB) ($Z=1$) versus vertical sleeve gastrectomy (VSG) ($Z=0$) bariatric surgery procedures on percent BMI change at 3 years post-surgery ($Y$). From a larger sample of 30,991 patients for which follow-up outcomes were partially observed, we selected $13,514$ patients who underwent RYGB or VSG between January 2005 and September 2015, with complete weight data at baseline (closest measurement pre-surgery, up to 6 months) and at follow-up (closest measurement within $\pm$ 90 days). See Table S.3 for a summary of the baseline characteristics of individuals who underwent bariatric surgeries.
@@
Subsequently, to impose MNAR, we additionally included a dependence on the outcome $Y$ by augmenting the fitted values $\widetilde{p}$ of $\textrm{pr}(R=1\mid Z,X)$ via: $\widetilde{p}\ \mapsto p$ where $p=\textrm{expit}\{\textrm{logit}(\widetilde{p})+Y^{*}[2-Z+0.3(\text{gender})+0.3(\text{race})-(\text{diabetes})-0.4(\text{statin})-0.5(\text{insulin})]\}$ where ${Y^{*}}$ is the standardized 3-year BMI change. Finally, we trimmed $p$ using $0.05$ and $0.95$. To generate MNAR outcomes on the sample of 13,514 patients, we sampled $R$ according to a Bernoulli with probability given by the trimmed $p$. The resulting marginal missingness probability was 30%. We considered double-sampling subsamples of size 1,000 and 1,500 from all patients with initially missing outcomes using simple random sampling, and estimated the double-sampling probability based on the sample proportion. For each of the two resulting data sets, we computed the QTEs at a sequence of quantile levels (i.e., $0.1,0.3,0.5,0.7,0.9$), and compared the point estimates and 95% pointwise bootstrap CIs using the proposed method in Section 3. A benchmark QTE estimate for each quantile level was estimated using the full data set (before artificial imposition of missing outcomes). For comparison, we also estimated QTEs using the complete-case data set ($R=1$). We computed 95% confidence bands using the data set with 1,500 double-sampled patients and compared to the pointwise bootstrap confidence intervals. Additionally, as the effects of the two bariatric surgeries may be different among patient with diabetes or dyslipidemia diagnosis, we repeated the above analyzes stratifying by diabetes and dyslipidemia diagnosis categories.
In Figure 3, the left column displays the pointwise QTE estimates and 95% CIs across quantile levels , and the right column compares pointwise CIs to confidence bands constructed in Section 3.6. Panels (A) to (C) represent different data subsets: (A) for all 13,514 patients, and (B) and (C) are stratified by diabetes status and dyslipidemia diagnosis, respectively. In panel (A), the confidence band is slightly wider than the pointwise CIs, likely attributable to idiosyncrasies in the real data application. The specific estimated QTE values and 95% pointwise CIs are listed in Table S.4 in the Supplementary Materials. The results show that complete case estimates are biased compared to the benchmark estimates using the full data, especially at low quantile levels. The proposed IPW-WTQE estimators were unbiased, with the setting with 1,500 subjects in the second-stage sample showed greater efficiency (i.e. with narrower 95% CIs) compared to the setting with the smaller subsample. Heterogeneous QTEs were observed across quantile levels as well as at different strata: when compared to VSG, RYGB appears to be associated with greater percent BMI loss at 3 years: (i) at low tails of the percent BMI change distribution, and (ii) among patients with diabetes or dyslipidemia diagnosis at most of the selected quantile levels.
@@ 6 Discussion
The work of this paper is motivated, in part, by Koffman et al. (2021), who report on an NIH-funded study that used a telephone survey to investigate missingness in EHR-based weight measurements, relevant to a broader study of long-term outcomes following bariatric surgery. Key to that publication is that it demonstrated the potential for the use of double-sampling in practical settings, specifically as a means to collect information on covariates that are potentially MNAR and learn about missingness mechanisms. The contribution of this paper, then, is to propose a novel approach for estimation and inference for causal weighted QTEs, one that combines the readily-available data in, say, the EHR, with additional data collected on a sub-sample of study units with otherwise missing outc ome information, and a set of proposed identifying assumptions. When estimating contrasts involving the mean of the counterfactual distribution such as the ATE, one may employ $g$-methods that first compute a conditional ATE, conditional on $X$, and then marginalize to recover an estimate of the population causal effect (Hernán and Robins, 2020). However, due to Jensen’s inequality, integrating a conditional QTE will not yield the population QTE (Sun et al., 2021). Motivated by this, we propose a novel inverse-probability weighted estimator as well as a bootstrap-based approach for valid inference, either pointwise for a single quantile or uniformly for a range. The procedure involves first estimating propensity scores and double-sampling probabilities, and then plugging these into an estimating function for the weighted QTE. In a simulation study, we consider settings where the QTEs are either homogeneous or heterogeneous across quantiles, verify theoretical estimation properties in a finite sample setting and demonstrate validity of the proposed bootstrap-based confidence intervals/bands.
@@
Key to the proposed methods is that no assumptions are required regarding the missingness mechanism in the original data. This, arguably, is a critical benefit because mechanisms underlying $R$, especially in the increasingly common setting of using EHR data, are typically outside of the control of the research team and potentially quite complex. Thought experiments on assumptions necessary for progress with the original data (e.g., to consider the potential validity of MAR) can be challenging and inconclusive. Thus, one way of framing the proposed methods is that assumptions regarding $R$ (as would have to be the case if only the original data were available) are replaced with the assumptions regarding the double-sampling indicator $S$; that is, Assumptions 3.2 and 3.2. In practice, depending on the context and the nature of the additional data to be collected, consideration of these assumptions may be fairly straightforward. For example, suppose the additional data collection involves manual chart review or natural language processing of a sub-sample chosen by the research team (Weiskopf et al., 2019), Assumptions 3.2 and 3.2 may hold trivially by design, since no involvement of the study subject is needed. In other settings, however, in particular when study subject participation is needed (e.g., to respond to a survey or phone call), assessment of these assumptions may require more detailed consideration. However, even in these situations, researchers may pursue additional design options that may shed light. For example, they may employ intensive, repeated efforts to contact study subjects and collect the data and/or engage with study subjects who have complete information in the original data to learn about potential recall bias (Haneuse et al., 2016; Koffman et al., 2021). EHR studies often encompass vast datasets, yet the ability to reach out to patients typically does not grow with sample size. Given that the double-sampling framework can be strategically planned from the beginning, it presents an opportunity to allocate resources effectively within the budgetary confines. Our ongoing research aims to find an optimal sampling rule that minimizes the semiparametric efficiency bound subject to a budget constraint.
@@
Acknowledgements
@@
Supplementary Materials
Web Appendices S.1, S.2, and S.3 referenced in Section 3, Web Appendices S.4, Tables S.1 to S.2, and Figures S.1 to S.8 referenced in Section 4, Web Appendices S.5 and Tables S.3, S.4 referenced in Section 5, and the R code referenced in Section 4 and 5 are available with this paper at the Biometrics website on Oxford Academic.
Data Availability
The data that support the findings in this paper cannot be shared publicly due to constraints imposed by Data Use Agreements with Kaiser Permanente.
Figure 1: Pointwise relative bias and 95% confidence interval coverages of QTE estimates across quantile levels. Row (A): Homogeneous QTE setting. Row (B): Heterogeneous QTE setting. Left and middle columns: Relative bias. Right column: Pointwise 95% confidence interval coverage Estimators (I)–(V) denote the five estimators in Section 4.2. The asymptotic variance and bootstrap variance are defined in Section 4.2. Total 10,000 datasets, each with sample size 10,000. Abbreviation: VAR, variance.
Figure 2: Comparison of pointwise 95% confidence intervals versus uniform bands for the point QTE estimates of a realization. The left panel is in homogeneous QTE setting. The right panel is in heterogeneous QTE setting. Pointwise confidence interval were estimated using the asymptotic and bootstrap variance, as defined in Section 4.2; the confidence band is outlined in the same section. Abbreviation: CI, confidence interval.
Figure 3: (A) The $n$=13,514 sample. (B) Stratified by diabetes status. (C) Stratified by dyslipidemia diagnosis. Left column: Pointwise QTE estimates and 95% pointwise bootstrap CIs across quantile levels; point estimates are marked with dots. Right column: Comparison of 95% pointwise boostrap CIs versus uniform confidence bands; CIs are the dashed lines and CBs are the shaded areas. Abbreviation: DS, double-sample; CI, confidence interval; CB, confidence band.